Question 1: You are waiting to board at an airport for a trip to Korea. You decided to buy cigarettes and Swiss chocolates for your friends at a shop in the airport. You have $30 only. A price for one pack of cigarettes is $2 and a price for one bar of Swiss chocolates is $3. You also learned from a travellers book that you can only bring 10 packs of cigarettes into the country at the maximum and are not allowed to buy more than that (cigarettes rationing) 1. Write down the budget set and budget line in equations for the case without regulation on cigarettes amount. Draw a budget line for each of two cases in a same graph: without the regulation on cigarettes amount and with the regulation in place. Put Swiss chocolates on the y axis and cigarettes on the x axis. Label x and y intercepts. 2. 3. Suppose the price of a bar of Swiss chocolates increases to $4 due to an increase in cocoa powder. Draw a new budget line with the cigarettes rationing under the new price in the same graph. 4. Can you afford an bundle (cigarettes-7 packs, chocolates-5 bars) under the old price of chocolates (i.e.p2 3 )? How about under the new price of chocolates (i.e.p2 4)?
Answer #1

(A) - If we dont keep any regulation on the amount of cigg. then by keeping cigg. on the X axis and swiss chocolates on the Y axis we get the following budget lines and budget sets-

15x + 0y = 30 (15 cigg)

9x + 4y = 30 (9 cigg , 4 swiss chocolate)

3x + 8y = 30 (3 cigg, 8 chocolates)

0x + 10y =30(10 chocolates)

12x + 2y = 30 (12 cigg, 2 chocolates)

6x + 6y = 30 (6 cigg, 6 chocolates)

The price for 1 pack of cigg is $3 (x) and the price of one pack of swiss chocolate is $ 2.(y)

(B) The budget line with the equation 12x + 2y = 30 represents the line with no regulation on the amount of ciggerettes.

The budget line with the equation 9x + 4y = 30 represents the line with the regulation on the amount of ciggerettes.

This is to state that x represents the price of one pack of ciggerette i.e $2 and y represents the price of each swiss chocolate i.e $ 3.

The lines have been depicted in the diagram below-

ID 0 O 2 36 1 8 lo 1 12 Nlo

C- In the above diagram the budget line with the equation 5x + 5y = 30 reresents the condition when the price of the swiss chocolate has been raise to $4. In such a case with the regulation still aplicable on the amount of cigg. we can purchase 5 units of pack of ciggerttes and 5 swiss chocolates in a given budget of $30.

D-   Yes we can buy 7 pack of cigg. and 5 swiss chocolates under the price of the swiss chocolates being being $3 er chocolate. This is because 7 cigg. will cost $ 14 and 5 swiss chocolates will cost$ 15 which amounts to total of $29 that is under our budget of $ 30.

No we cannot buy 7 pack of cigg. and 5 swiss chocolates when the price of the chocolates rises to $4 per chocolate because $ 14 + $ 20 = $ 34 which goes beyond our budget of $ 30.